OpenStudy (anonymous):

Byron is 3 years older than Doug. The product of their ages are 40. How old is Doug?

OpenStudy (anonymous):

Let the age of Doug be x then Byron is (x+3) years old Now, According to the question, x*(x+3)=40

OpenStudy (anonymous):

Can u solve for x

OpenStudy (anonymous):

Well then I cannot

OpenStudy (anonymous):

U will get a quadratic equation........U know how to solve it

OpenStudy (anonymous):

Yes thank you

OpenStudy (anonymous):

x(x+3)=40 x^2 + 3x=40 x^2+3x-40=0 x^2 +8x -5x -40=0 x(x+8) -5(x+8)=0 (x-5)(x+8)=0

OpenStudy (anonymous):

Now, can u ?

OpenStudy (anonymous):

Yes thanks

OpenStudy (anonymous):

Welcome